OpenStudy (anonymous):

log 15 ( x 2 + 13) = log 15 ( −9 x − 1)

OpenStudy (anonymous):

log 15 divides out leaving\[(13+2 x)=(-1-9 x) \]Solve for x.\[(13+2 x)-(-1-9 x)=0 \]\[14+11 x=0\]\[x=-\frac{14}{11} \]